OpenStudy (anonymous):

You have 15 pennies in your pocket. Of those pennies, 3 are Canadian. Suppose you pick a penny out of your pocket at random. Find P(not Canadian). A.5 B.1/5 C.6/5 D.4/5

sam (.sam.):

3 are Canadian 15-3= 12 not Canadian 12/15

sam (.sam.):

or 4/5

OpenStudy (anonymous):

15 * 3 = 45

OpenStudy (anonymous):

so the answer is 4/5.

OpenStudy (anonymous):

P ( Canadian) = 3/15 = 1/5 -> P (not canadian) = 1 - 1/5 = 4/5